OpenStudy (anonymous):

does anyone know this one ? 2(-6-4)2 - 8.3 the 2 is to the second power

OpenStudy (anonymous):

200-24=176

OpenStudy (anonymous):

176..

OpenStudy (anonymous):

2(-10)^2 - 24 =2(100) - 24 =200 - 24 = 176

OpenStudy (anonymous):

176

OpenStudy (anonymous):

What is the Area of a rectangular room that has a length of 7 feet and a width of 5 fee

OpenStudy (anonymous):

How much money is 15% of $20?

OpenStudy (anonymous):

Perhaps you should post 1 question at a time so you learn how to do the problem...